OpenStudy (anonymous):

Another one... Find the Cartesian inequation for the region represented by | z+5+9 i | < (1/6) | z+7+3 i | ... Could anyone please help me with this problem?

OpenStudy (anonymous):

simply use z=x+iy arrange real and imaginary parts use mod(a+ib)=sqrt(a^2+b^2) u will get the required equation